The unit of measure for a line is in inches. Use the distance formula to find the length of the line with points at (-3,0) and (4,3).
Answer by palanisamy(496)

Length = SQRT [ (-3-4)^2+(0-3)^2]=SQRT[ 49+9] =SQRT(58)